OpenStudy (anonymous):

The function S(n) is defines below. Which of the following are equal to S(10)?Check all that apply S(n)=5n-3 A)S(9)+5 B)53 C)47 D)S(7)+S(3)

OpenStudy (anonymous):

S(10)=5*10-3=50-3=47 so C is right

OpenStudy (anonymous):

so its only c?

OpenStudy (anonymous):

lets check them

OpenStudy (anonymous):

S(9)+5=5*9-3+5=45-3+5=47 so A is right too

OpenStudy (anonymous):

now try for S(7)+S(3) what will u get?

OpenStudy (anonymous):

S(10)

OpenStudy (anonymous):

so no?

OpenStudy (anonymous):

Only A and C

OpenStudy (anonymous):

what is the value of S(7)+S(3) ?

OpenStudy (anonymous):

s10?

OpenStudy (anonymous):

no S(7)=5*7-3=32 S(3)=5*3-3=12 then S(7)+S(3)=12+32=43 so Only A and C

OpenStudy (anonymous):

ok thank you